5x-9y=-17
5x+33= y
solve and find the ordered pair

Answer by ewatrrr(24785)

 
Hi,
5x-9y=-17
5x+33= y
5x-9y = -17
5x -y = -33
5x -9y = -17
-5x +y = 33
-8y = 16
y = -2 and x = -35/5 = -7
Ordered pair (-7,-2) the solution for this system of equations